# Know Unreported Roadway Incidents in Real-time: A Deep Learning Framework for Early Traffic Anomaly Detection

### v1

@@ Abstract
The goal of this research is to know traffic anomalies as early as possible. A traffic anomaly refers to a generic incident on the road that influences traffic flow and calls for urgent traffic management measures. “Knowing” the occurrence of a traffic anomaly is twofold: the ability to detect this anomaly before it is reported anywhere, or it may be such that an anomaly can be predicted before it actually occurs on the road (e.g., non-recurrent traffic breakdown). Either way, the objective is to inform traffic operators of unreported incidents in real time (and as early as possible), regardless of whether it is being reported later or not. The key is to stay ahead of the curve. Time is of the essence.
Conventional automatic incident detection (AID) has relied heavily on all incident reports exclusively for training and evaluation. However, these reports suffer from a number of issues, such as delayed reports, inaccurate descriptions, false alarms, missing reports, and incidents that do not necessarily influence traffic. Relying on these reports to train or calibrate AID models hinders their ability to detect traffic anomalies effectively and timely, even leading to convergence issues in the model training process. Moreover, conventional AID models are not inherently designed to capture the early indicators of any generic incidents. It remains unclear how far ahead an AID model can report incidents. The AID applications in the literature are also spatially limited because the data used by most models is often limited to specific test road segments. To solve these problems, we propose a deep learning framework utilizing prior domain knowledge and model-designing strategies. This allows the model to detect a broader range of anomalies, not only incidents that significantly influence traffic flow but also early characteristics of incidents along with historically unreported anomalies. We specially design the model to target the early-stage detection/prediction of an incident. Additionally, unlike most conventional AID studies, we use widely available data, enhancing our method’s scalability. The experimental results across numerous road segments on different maps demonstrate that our model leads to more effective and early anomaly detection. Our framework does not focus on stacking or tweaking various deep learning models; instead, it focuses on model design and training strategies to improve early detection performance.
@@ 1 Introduction
Traffic congestion grievously disrupts the everyday lives of urban residents and causes substantial economic losses. It can be categorized into recurring and non-recurring congestion. Recurring congestion occurs periodically when traffic volume exceeds the road’s capacity. In contrast, non-recurring congestion, also known as traffic anomalies, is caused by incidents such as crashes, work zones, and special events. In the U.S., non-recurring congestion accounts for over half of the total congestion (Federal Highway Administration (FHWA), 2020b). Mitigating non-recurring impacts requires accurately knowing roadway incidents in advance and proactive management strategies as opposed to being reactive. Figure 1 illustrates the FHWA timeline for managing non-recurring events, highlighting that the effectiveness of proactive strategies heavily relies on the timely detection and verification of incidents. However, incident reports in the real world are typically delayed and subject to multiple layers of verification, not to mention a large number of unreported incidents. Acknowledging that incident reports are oftentimes late or missing, traffic operators may find their actions too late by the time they are informed of non-recurrent congestion. To assist traffic operators in implementing real-time control measures (F. Ahmed and Hawas, 2015; Sheu, 2002; Z. Ke, Zou, et al., 2024) and travelers in planning their journeys (Z. Ke, Duan, and Qian, 2024; Xie et al., 2020), it is crucial to verify incident reports more promptly, or better yet, detect incidents before they are reported and effectively identify unreported incidents. Our goal is to alert traffic operators of an anomaly before $T_{1}$, the current practice, by as much as we can, and ideally even before $T_{0}$ referred in Figure 1.
@@
Conventional automatic incident detection (AID) methods may assist in verifying incident reports, but it is unclear if AID is able to detect incidents earlier than the reports themselves after the fact and efficiently identify unreported incidents. This is because they are typically tuned/trained and evaluated based solely on all incident reports, thereby inheriting the same delays and missing issues. These issues will be further discussed in detail later in this section. This paper aims to overcome the bottlenecks caused by incident reports in conventional AID methods and management strategies, with the goal of detecting or predicting anomalies as early as possible. Here, we use “early anomaly detection” to indicate that our problem is fundamentally different from conventional AID and from network traffic speed/counts prediction. An anomaly in this paper is defined not only as a reported incident but also as any incident that may not be reported at the time, potentially reported later, or not reported at all, yet significantly impacts traffic flow in the near term. Efficiently detecting all these anomalies, especially unreported incidents and incidents that have not yet been reported, is essential for the timely implementation of proactive control measures and the overall incident management process.
Given the exact anomaly occurrence time (namely $T_{0}$ referred in Figure 1) is usually unknown in practice, we do not necessarily differentiate “detecting” an ongoing but unreported anomaly and “predicting” an anomaly that has not occurred yet. The goal is to know an unreported anomaly in advance, regardless of whether it has occurred. It is critical to clarify we do not intend to “predict” the occurrence of an incident before it actually occurs, since generally incident occurrence is rare and random. Our intention is to know the anomaly as early as we can, even if there is very little sign of an incident yet. This means that at the time of anomaly detection, it is likely an incident has occurred, but not reported anywhere (including Waze and social media, which is considered more timely than other traditional manners). It is also possible (but unlikely) an incident may not have occurred, but a certain traffic/weather condition may lead to high risks resulting its occurrence in the near future. Again, our algorithm does not differentiate this “detection” or “prediction” nature, thus lending its name to “early anomaly detection”. We also do not intend to detect or predict the attributes of an incident, its type, duration, etc.
@@
The issue of delayed incident report arises from the long reaction time between the incident occurrence and when they are reported (Gu et al., 2016; ElSahly and Abdelfatah, 2022). Figure 2 illustrates the delay of incident reports by displaying the relationship between Waze reports and speed on an I-70 highway segment in Howard County, MD. The blue line represents the speed on the day of the incident report, while the other dotted lines represent the speed on the same days of the week and at the same time, providing a recurrent reference. It can be seen from the figure that Waze incident reports are significantly delayed compared to the impact on speed. As a result of overlooking delays in incident reports, those features prior to incident reports are often missing from training an AID model. Due to the lack of training samples in the early stage of an incident, the model is unable to learn from these features signing an upcoming or ongoing incident, hindering timely detection. It is worth noting that many studies emphasize achieving near-zero mean time to detection (MTTD) as evidence of timely detection:
@@
where $N_{\text{detected}}$ refers to the number of incident report being detected, $t_{i}^{\text{report}}$ refers to the start time of the report, and $t_{i}^{\text{alarm}}$ refers to the time when the model begins to trigger an alert. However, this is because their evaluation process compares detection times to delayed incident reports, so a near-zero MTTD simply indicates that the alarm is triggered shortly after the incident is reported, not when it actually occurs.
@@
In addition, the issue of missing incident report is mainly because incident reports still largely rely on either fixed location sensors that are extremely limited or passing drivers to report actively (e.g. Waze, X, etc.). In fact, many incidents or traffic anomalies go unreported (Turner et al., 2015). Figure 3 illustrates examples of significant anomalies that were not reported on the same segment as in Figure 2. The blue line highlights anomalies without corresponding incident reports during those times, while the other dashed lines represent the speed on the same days of the week and at the same time, providing a recurrent reference. Significant anomalies can be observed even without the presence of an incident report. Those data under unreported anomalies are otherwise not included in the incident detection/prediction models. Even worse, they are labeled as data under recurrent traffic conditions, which could mislead the training of AID models.
@@
In summary, overlooking the issues related to incident reports leads existing AID methods to fit their results to the reports themselves (which may be delayed, false, or missing) rather than to true and timely anomalies, particularly the early stages of an incident or unreported incidents, limiting conventional AID models’ performance.
@@
Besides performance issues, another key challenge for conventional AID methods today is their scalability and generalizability. Most of the data used in AID methods, such as loop detector data, traffic camera data, and high-precision GPS data, are not widely available or can be very costly in real-time. For instance, real-time 24/7 loop detector data is limited to a very small number of segments or none in most regions. Social media data used by a few studies have also faced more scraping restrictions nowadays. On the other hand, data sources such as INRIX (www.inrix.com), HERE(www.here.com), and TomTom (www.tomtom.com) provide probe vehicle speed data, while Waze (www.waze.com) and State DOT offer incident reports. These data sources cover almost all major roads, providing the foundation for large-scale generalization of AID methods. The more ubiquitous data coverage of spaces and time enables a better understanding of network traffic flow and, thus, early anomaly detection on most road segments. It is essential to develop an incident detection/prediction model relying on those data so that we can afford to detect/predict incidents virtually anywhere and anytime. However, very few models are built exclusively on such data.
@@
To address the above problems, we will propose an early anomaly detection system that can learn and capture a wider range of anomalies, i.e., not only fits incident reports but also detects early incident features and unreported anomalies, enhancing the model’s practical significance. In addition, to ensure scalability and generalizability, the inputs to the new model should be limited to widely available low-cost data rather than data from extremely selected locations. Our contributions can be summarized as:
@@
  • We propose an early anomaly detection model employing several training and model design strategies. Our model not only aligns well with true incident reports but also performs early detection or prediction of incidents before they are reported later after the fact or not reported at all. This model is fundamentally different from conventional AID models.
@@
  • In contrast to most datasets in the literature, our method does not require manual selection of anomaly reports during training, saving a significant amount of manpower and increasing its scalability.
@@
  • Our model uses ubiquitously available and low-cost data, increasing its scalability and generalizability. Unlike most existing studies that test one or two specifically selected segments, we conducted experiments on ten segments across two road networks, which can be generalized to any size of road segments in the network.
@@ 2 Literature Review
Comparative methods collect traffic data such as flow, speed, and occupancy through detectors and compare them with past incident patterns. Alarms are triggered when the collected data matches a pre-defined incident pattern. Representative comparative methods include the California (Payne and Tignor, 1978), Manchester (Persaud and Hall, 1989), and Minnesota (Stephanedes and Chassiakos, 1993) algorithms. These methods are highly inspiring and have pioneered the analysis and extraction of incident characteristics. However, they largely remain at the level of manually summarizing incident characteristics, relying on pre-defined simple structures or patterns to determine whether to trigger an alert. The limited data mining capacity results in a high false alarm rate and mean time to detection.
Due to the richness and complexity of traffic data, researchers have begun to consider using artificial intelligence for data mining to detect incidents. Cheu and Ritchie (1995) used multi-layer feedforward neural networks (MLF) for incident detection, outperforming traditional methods on simulated data and small field data containing 9 incidents. To further validate feasibility on real-world data, Dia and Rose (1997) used data collected from Tullamarine Highway, Melbourne, demonstrating their feasibility on real datasets. Abdulhai and Ritchie (1999) optimized the model, proposing a Bayesian-based probabilistic neural network (BPNN), and trained the model using data from I-880, California, and I-35W, Minnesota. Jin et al. (2002) further optimized the model by proposing the constructive probabilistic neural network (CPNN) and conducted studies on the same experimental section of I-880. Given the strong performance of support vector machines (SVM) on binary classification problems, Yuan and Cheu (2003) proposed an SVM-based model and validated it on the datasets used by Jin et al. (2002) and Abdulhai and Ritchie (1999). To address the lack of incident samples, Lin et al. (2020) used generative adversarial networks (GANs) to generate more incident samples and use them combined with real incident samples to train an SVM classifier. Similarly, Li et al. (2022) used GANs on the same dataset as Lin et al. (2020) but replaced the SVM classifier with a temporal and spatially stacked autoencoder (TSSAE).
@@
In addition to input data, another long-neglected issue is anomaly labels. Anomaly labels typically come from incident reports, but due to false reports and insignificant incidents, model training often fails to converge. Therefore, it has long been necessary to manually filter out real and impactful accidents. For example, Dia and Rose (1997) manually selected 60 out of 385 accidents, Abdulhai and Ritchie (1999) manually filtered and organized accident reports on I-880 California and I-35 Minnesota, and Chakraborty et al. (2019) combined manual selection with traffic cameras to filter incident reports. Subsequent researchers have widely used these datasets. Filtering and organizing incident reports requires significant manual effort, limiting the model’s generalizability. We also cannot utilize the existing abundance of incident reports to achieve model generalization. Furthermore, even after complex manual filtering, incident reports still have inherent issues, such as delays and missing, leading to a lack of relevant anomaly samples in the training data, which restricts the model’s performance. To address this issue, Petty et al. (1996) used driver manually calibrated data, but this is obviously impractical for most datasets. Current research still aligns prediction results with incident reports rather than anomalies. We need a model capable of achieving broader anomaly detection.

### v2

# Know Unreported Roadway Incidents in Real-time: Early Traffic Anomaly Detection

@@ Abstract
This research aims to know traffic anomalies as early as possible. A traffic anomaly refers to a generic incident on the road that influences traffic flow and calls for urgent traffic management measures. ‘Knowing” the occurrence of a traffic anomaly is twofold: the ability to detect this anomaly before it is reported anywhere, or it may be such that an anomaly can be predicted before it actually occurs on the road (e.g., non-recurrent traffic breakdown). In either way, the objective is to inform traffic operators of unreported incidents in real time and as early as possible. The key is to stay ahead of the curve. Time is of the essence.
Conventional automatic incident detection (AID) methods often struggle with early detection due to their limited consideration of spatial effects and early-stage characteristics. Therefore, we propose a deep learning framework utilizing prior domain knowledge and model-designing strategies. This allows the model to detect a broader range of anomalies, not only incidents that significantly influence traffic flow but also early characteristics of incidents along with historically unreported anomalies. We specially design the model to target the early-stage detection/prediction of an incident. Additionally, unlike most conventional AID studies, our method is highly scalable and generalizable, as it is fully automated with no manual selection of historical reports required, relies solely on widely available low-cost data, and requires no additional detectors. The experimental results across numerous road segments on different maps demonstrate that our model leads to more effective and early anomaly detection.
@@ 1 Introduction
Traffic congestion grievously disrupts the everyday lives of urban residents and causes substantial economic losses. It can be categorized into recurring and non-recurring congestion. Recurring congestion occurs periodically when traffic volume exceeds the road’s capacity. In contrast, non-recurring congestion, also known as traffic anomalies, is caused by incidents such as crashes, work zones, and special events. In the U.S., non-recurring congestion accounts for over half of the total congestion (Federal Highway Administration (FHWA), 2020b). Mitigating non-recurring impacts requires accurately knowing roadway incidents in advance and proactive management strategies as opposed to being reactive. Figure 1 illustrates the FHWA timeline for managing non-recurring events, highlighting that the effectiveness of proactive strategies heavily relies on the timely detection and verification of incidents. However, incident reports in the real world are typically delayed and subject to multiple layers of verification, not to mention a large number of unreported incidents. Acknowledging that incident reports are oftentimes late or missing, traffic operators may find their actions too late by the time they are informed of non-recurrent congestion. To assist traffic operators in implementing real-time control measures (F. Ahmed and Hawas, 2015; Sheu, 2002; Z. Ke, Zou, et al., 2024) and travelers in planning their journeys (Z. Ke, Duan, and Qian, 2024; Xie et al., 2020), it is crucial to verify incident reports more promptly, or better yet, detect incidents before they are reported and effectively identify unreported incidents. Our goal is to alert traffic operators of an anomaly before $T_{1}$, the current practice, by as much as we can, and ideally even before $T_{0}$ referred in Figure 1 in situations where the risk of roadway incidents or non-recurrent traffic breakdown is extremely high.
@@
Conventional automatic incident detection (AID) methods may assist in verifying incident reports, but it is unclear if AID is able to detect incidents earlier than the reports themselves after the fact and efficiently identify unreported incidents. This is because they are typically tuned/trained and evaluated based solely on incident reports, thereby inheriting the same delays and missing issues. These issues will be further discussed in detail later in this section. This paper aims to overcome the bottlenecks caused by incident reports in conventional AID methods and management strategies, with the goal of detecting or predicting anomalies as early as possible. Here, we use “early anomaly detection” to indicate that our problem is fundamentally different from conventional AID and from network traffic speed/counts prediction. An anomaly in this paper is defined not only as a reported incident but also as any incident that may not be reported at the time, potentially reported later, or not reported at all, yet significantly impacts traffic flow in the near term. Efficiently detecting all these anomalies, especially unreported incidents and incidents that have not yet been reported, is essential for the timely implementation of proactive control measures and the overall incident management process.
Given the exact anomaly occurrence time (namely $T_{0}$ referred in Figure 1) is usually unknown in practice, we do not necessarily differentiate “detecting” an ongoing but unreported anomaly and “predicting” an anomaly that has not occurred yet. The goal is to know an unreported anomaly in advance, regardless of whether it has occurred. It is critical to clarify we do not intend to “predict” the occurrence of an incident before it actually occurs, since generally the incident occurrence is rare and random. Our intention is to know the anomaly as early as we can, even if there is very little sign of an incident yet. This means that at the time of anomaly detection, it is likely that an incident has occurred but not reported anywhere (including Waze and social media, which is considered more timely than other traditional manners). It is also possible (but unlikely) that an incident may not have occurred, but a certain traffic/weather condition may lead to high risks resulting in its occurrence in the near future. Again, our algorithm does not differentiate this “detection” or “prediction” nature, thus lending its name to “early anomaly detection”. We also do not intend to detect or predict the attributes of an incident, its type, duration, etc.
@@
The issue of delayed incident reports arises from the long reaction time between the incident occurrence and when they are reported (Gu et al., 2016; ElSahly and Abdelfatah, 2022; Coursey et al., 2024). Figure 2 illustrates the delay of incident reports by displaying the relationship between Waze reports and speed on an I-70E highway segment in Howard County, MD. The blue line represents the speed on the day of the incident report, while the yellow and green dotted lines represent the speed on the same days of the week and at the same time, providing a recurrent reference. It can be seen from the figure that Waze incident reports are significantly delayed compared to the impact on speed. As a result of overlooking delays in incident reports, those features prior to incident reports are often missing from training an AID model. Due to the lack of training samples in the early stage of an incident, the model is unable to learn from these features signing an upcoming or ongoing incident, hindering timely detection. It is worth noting that many studies emphasize achieving near-zero mean time to detection (MTTD) as evidence of timely detection:
@@
where $N_{\text{detected}}$ refers to the number of incident reports being detected, $t_{i}^{\text{report}}$ refers to the start time of the report, and $t_{i}^{\text{alarm}}$ refers to the time when the model begins to trigger an alert. However, this is because their evaluation process compares detection times to delayed incident reports, so a near-zero MTTD simply indicates that the alarm is triggered shortly after the incident is reported, not when it actually occurs.
@@
In addition, the issue of missing incident reports is mainly because incident reports still largely rely on either fixed location sensors that are extremely limited or passing drivers to report actively (e.g., Waze, X, etc.). In fact, many incidents or traffic anomalies go unreported (Turner et al., 2015). Figure 3 illustrates examples of significant anomalies that were not reported on the same segment as in Figure 2. Significant anomalies can be observed even without the presence of an incident report. Those data under unreported anomalies are otherwise not included in the incident detection/prediction models. Even worse, they are labeled as data under recurrent traffic conditions, which could mislead the training of AID models.
@@
In summary, overlooking the issues related to incident reports leads existing AID methods to fit their results to the reports themselves (which may be delayed, false, or missing) rather than to true and timely anomalies, particularly the early stages of an anomaly or unreported anomalies, which potentially are incidents, limiting conventional AID models’ performance.
@@
Besides performance issues, another key challenge for conventional AID methods today is their scalability and generalizability. Most of the data used in AID methods, such as loop detector data, traffic camera data, and high-precision GPS or radar sensor data, are not widely available or can be very costly in real-time. For instance, real-time 24/7 loop detector data is limited to a very small number of segments or none in most regions, let alone high-precision radar sensor data. Social media data used by a few studies have also faced more scraping restrictions nowadays. On the other hand, data sources such as INRIX (www.inrix.com), HERE(www.here.com), and TomTom (www.tomtom.com) provide probe vehicle speed data, while Waze (www.waze.com) and State DOT offer incident reports. These data sources cover almost all major roads, providing the foundation for large-scale generalization of AID methods. The more ubiquitous data coverage of spaces and time enables a better understanding of network traffic flow and, thus, early anomaly detection on most road segments. It is essential to develop an incident detection/prediction model relying on those data so that we can afford to detect/predict incidents virtually anywhere and anytime. However, very few models are built exclusively on such data.
@@
To address the above problems, we will propose an early anomaly detection system that can learn and capture a wider range of anomalies, i.e., not only fits incident reports but also detects early incident features and unreported anomalies, especially those affecting broader areas. In addition, to ensure scalability and generalizability, the inputs to the new model should be limited to widely available low-cost data rather than data from extremely selected locations or manually selected data. Our contributions can be summarized as:
@@
  • We propose an anomaly detection model specifically designed for early-stage incident detection and prediction. Our model not only aligns well with true incident reports but also detects or predicts incidents before delayed reporting or in cases where incidents remain unreported. This model is fundamentally different from conventional AID models.
@@
  • In contrast to most methods in the literature, the entire process is designed in an automated framework, which does not require manual selection of anomaly reports during training, thereby enhancing scalability and efficiency.
@@
  • Our model uses ubiquitously available and low-cost data with no additional detector required, increasing its scalability and generalizability. We conducted experiments on ten segments across two road networks, which can be generalized to any size of road segments in the network.
@@ 2 Literature Review
Comparative methods collect traffic data such as flow, speed, and occupancy through detectors and compare them with past incident patterns. Alarms are triggered when the collected data matches a pre-defined incident pattern. Representative comparative methods include the California (Payne and Tignor, 1978), Manchester (Persaud and Hall, 1989), and Minnesota (Stephanedes and Chassiakos, 1993) algorithms. These methods are highly inspiring and have pioneered the analysis and extraction of incident characteristics. However, they largely remain at the level of manually summarizing incident characteristics, relying on pre-defined simple structures or patterns to determine whether to trigger an alert. The limited data mining capacity results in a high false alarm rate and mean time to detection. Recent work, such as Grigorev et al. (2024), proposed comparing observed data with the monthly profile under similar traffic states. By computing the first-order derivative of the normalized Chebyshev distance, they designed a threshold-based method to capture significant temporal deviations from historical traffic patterns. Nevertheless, this method fundamentally depends on heuristic assumptions about incident characteristics.
Due to the richness and complexity of traffic data, researchers have begun to consider using artificial intelligence for data mining to detect incidents. Cheu and Ritchie (1995) used multi-layer feedforward neural networks (MLF) for incident detection, outperforming traditional methods on simulated data and small field data containing 9 incidents. To further validate feasibility on real-world data, Dia and Rose (1997) used data collected from Tullamarine Highway, Melbourne, demonstrating their feasibility on real datasets. Abdulhai and Ritchie (1999) optimized the model, proposing a Bayesian-based probabilistic neural network (BPNN), and trained the model using data from I-880, California, and I-35W, Minnesota. Jin et al. (2002) further optimized the model by proposing the constructive probabilistic neural network (CPNN) and conducted studies on the same experimental section of I-880. Given the strong performance of support vector machines (SVM) on binary classification problems, Yuan and Cheu (2003) proposed an SVM-based model and validated it on the datasets used by Jin et al. (2002) and Abdulhai and Ritchie (1999). To address the lack of incident samples, Lin et al. (2020) used generative adversarial networks (GANs) to generate more incident samples and use them combined with real incident samples to train an SVM classifier. Similarly, Li et al. (2022) used GANs on the same dataset as Lin et al. (2020) but replaced the SVM classifier with a temporal and spatially stacked autoencoder (TSSAE). He et al. (2023) proposed a reinforcement learning-based approach that assigns rewards based on temporal differences and demonstrated its effectiveness on synthetic data. However, the lack of real-world incident reports limited their analysis to only two real-world cases.
@@
In addition to input data, another long-neglected issue is anomaly labels. Anomaly labels typically come from incident reports. However, incident reports often contain numerous manual errors (Grigorev et al., 2024; He et al., 2023; Taheri Sarteshnizi et al., 2024), which can negatively impact prediction performance and even lead to model convergence issues. Therefore, it has long been necessary to manually filter out real and impactful incidents. Dia and Rose (1997) manually selected 60 out of 385 incidents, Abdulhai and Ritchie (1999) manually filtered and organized incident reports on I-880 California and I-35 Minnesota, Petty et al. (1996) used driver manually calibrated incident data, and Chakraborty et al. (2019) combined manual selection with traffic cameras to filter incident reports. Subsequent researchers have widely used these datasets or manually filtered their datasets (Zhu et al., 2021; Coursey et al., 2024). Filtering and organizing incident reports requires significant manual effort, limiting the model’s generalizability. To overcome this issue, Taheri Sarteshnizi et al. (2024) introduced Temporal Positioning of Flow-Density Samples (TP-FDS) derived from studying the flow-density relationship. They utilized loop-detector data at a 15-minute resolution and validated their approach through comparison with expert-labeled data. In order to avoid reliance on loop detector data and thereby enhance the generalizability, Grigorev et al. (2024) proposed an automatic method that filters incident reports based on the first-order derivative of temporal speed differences. Although the approach relies solely on speed data, its primary objective is to improve incident duration prediction. As a result, it primarily targets temporal anomalies and may struggle to identify spatial anomalies. Moreover, it does not address key challenges in traffic anomaly detection, such as detecting early-stage disruptions that fall just below predefined thresholds or accounting for missing incident reports. To the best of our knowledge, current speed-based automatic labeling methods generally fall short of systematically addressing diverse anomaly types, such as spatiotemporal anomalies, early-stage disruptions, and unreported incidents. Moreover, most current research still aligns prediction results with selected incident reports rather than anomalies. It is still unclear how newly generated labels can contribute to incident detection, especially in terms of timeliness. We need a model capable of achieving broader anomaly detection.